# Any function I can actually write down is measurable, right?

### v2

@@ Abstract.
is Lebesgue measurable. We also give similarly defined $g(x,y)$ with the property that the statement “$x\mapsto g(x,r)$ is measurable for every $r\in\mathbb{R}$” has large cardinal consistency strength (and in particular implies the consistency of $\mathsf{ZFC}$) and $h(m,x,y,z)$ such that $h(1,x,y,z),\allowbreak\dots,\allowbreak h(16,x,y,z)$ can consistently be the indicator functions of a Banach–Tarski paradoxical decomposition of the sphere.
@@ Introduction
While most students of mathematics encounter measure theory at some point, the basic machinery of measure theory (such as $\sigma$-algebras) is often regarded as fussy and irrelevant in practice. In particular, the issue of measurability of sets and functions seems to be something one only encounters in the first semester of graduate real analysis and never again. The common informal explanation of this is that any function one can ‘actually write down’ will be measurable and that one needs to go looking for trouble using the axiom of choice in order to find non-measurable objects. This is certainly a true statement empirically, [^1] It should be noted though that there are some fields, like the study of stochastic differential equations, in which $\sigma$-algebras and measurability are used more extensively as explicit bookkeeping machinery. and a fair amount of work has also been done by set theorists to explain the fact that one seems to ‘need the axiom of choice’ to build a non-measurable function or set. The two most notable results in this direction are probably Solovay’s famous construction of a model of $\mathsf{ZF}+\mathsf{DC}$ in which all subsets of $\mathbb{R}$ are measurable [29] and Shelah and Woodin’s work culminating in the thesis that ‘large cardinals imply that every reasonably definable set of reals is Lebesgue measurable’ [26]. Yet it has also been known since Gödel’s seminal construction of $L$ in the 1930s that the existence of fairly ‘explicit’ non-measurable objects is consistent with $\mathsf{ZFC}$. Despite how old these results are, it seems to me [^2] We will occasionally use the terms I, me, and my as convenient shorthand for ‘the author’ or ‘the author’s,’ as appropriate. that not many mathematicians are aware of just how explicit these explicit objects can be.
@@
Most of the notation we’ll use here is standard. To avoid collision with the notation for open intervals, we’ll use angle brackets for ordered tuples in expressions like $\langle 1,2\rangle$. We will also be using the logician’s convention of representing tuples of indexed variables with a bar. For instance, $\bar{x}$ might represent $\langle x_{1},x_{2},x_{3}\rangle$. In some places these notations will be used together. For instance, if $\bar{x}=\langle x_{1},x_{2},x_{3}\rangle$, then $\langle\bar{x},n\rangle$ is the same thing as $\langle x_{1},x_{2},x_{3},n\rangle$. We’ll be representing projection maps with $\pi$, usually with a subscript. For example, an expression like $\pi_{4}$ is meant to represent a projection from some product $X\times Y\times Z\times W$ down to $X\times Y\times Z$, but this will be spelled out explicitly when it is used. Finally, we are taking 00 to be a natural number and all suprema and infima are understood to be computed in the extended real numbers, which we will denote by $\overline{\mathbb{R}}$.
@@ 1. Independence results and Gödel’s $L$
With the majority of set-theoretic independence phenomena, the very basic core argument is the same. One exhibits models of something like $\mathsf{ZF}$ or $\mathsf{ZFC}$ both satisfying and failing to satisfy a given sentence. The primary difference is that models of $\mathsf{ZFC}$ are large and rich enough to encode the majority of mathematics. While models of set theory are obviously far harder to visualize than, say, groups with $6$ elements, sometimes the relative construction of these models can be visualized fairly clearly.
Like many mathematicians, set theorists’ intuition is often founded on cartoonishly simple mental pictures, such as the standard drawing of a model of $\mathsf{ZFC}$ (see Fig. 1). A model of $\mathsf{ZFC}$ is somewhat like a great tree, rooted at the empty set and with the ordinals forming a transfinitely tall trunk off of which the various sets coding familiar mathematical objects branch. This picture immediately gives two coordinates which describe the broad structure of the majority of constructions of one model of $\mathsf{ZF}(\mathsf{C})$ from another: One can imagine a model extending another in height, leading to the idea of large cardinals, or one can imagine a model that extends another in width, as is the case in set-theoretic forcing, or a model which is the result of pruning another to some narrower width, which is the very basic idea of inner models such as Gödel’s $L$. Many more elaborate proofs involve performing several (or even infinitely many) of these constructions in succession, such as in Cohen’s original proof of the independence of choice from $\mathsf{ZF}$, which involves forcing to widen a model and then pruning down to a narrower symmetric submodel (a kind of inner model specialized for the context of forcing) to kill the axiom of choice by carefully cutting certain choice functions out of the model.
@@ 1.1. The easiest independence result
For any inaccessible $\kappa\in M$, $H_{\kappa}$ is a model of $\mathsf{ZFC}$ (see Fig. 2). This gives us what is probably the easiest $\mathsf{ZFC}$ independence result.
@@
Fix a model $M$ of $\mathsf{ZFC}$ with an inaccessible cardinal. Let $\kappa$ be the least inaccessible cardinal in $M$ (which exists by the fact that the cardinals are well-ordered). $H_{\kappa}^{M}$ is now a model of $\mathsf{ZFC}$ such that no set in $H_{\kappa}^{M}$ is itself a Grothendieck universe. Therefore $H_{\kappa}^{M}$ is a model of $\mathsf{ZFC}$ satisfying ‘there are no inaccessible cardinals.’ Since we now have a model of $\mathsf{ZFC}$ with an inaccessible cardinal and a model of $\mathsf{ZFC}$ without an inaccessible cardinal, the statement “there is an inaccessible cardinal” must be independent of $\mathsf{ZFC}$. ∎
@@ 1.2. The narrowest inner model
Given a model $M$ of $\mathsf{ZF}$, a class $N\subseteq M$ is an inner model if $N$ contains all of the ordinals in $M$, $N$ is transitive (i.e., satisfies that if $x\in N$ and $y\in x$, then $y\in N$), and $N$ is a model of $\mathsf{ZF}$.
Gödel showed how to build the ‘narrowest’ inner model $L$, which can be thought of as the model of $\mathsf{ZF}$ ‘generated’ by the ordinals (see Fig. 3), although this description is hard to make precise in a robust way. When it’s important to indicate that we are thinking about $L$ inside a particular model, we typically denote this with a superscript (i.e., $L^{M}$ is $L$ built in $M$).
@@ 1.3. The definable well-ordering of $L$
One particularly important aspect of the fine structure of $L$ is that there is a precise mechanism that results in the axiom of choice holding, unlike in arbitrary models of $\mathsf{ZFC}$. One way to understand this mechanism is that in the construction of $L$, sets are added ‘carefully’ so that it is possible to assign an ordinal ‘serial number’ to each set. In other words, there is a bijection $f:\operatorname{Ord}\to L$ that is definable in $L$. This induces a definable global well-ordering $<_{L}$ of $L$, which then gives a definable global choice function: Whenever you need to pick an element of a set $X$, you can just take the $<_{L}$-least element of $X$ (which always exists since $<_{L}$ is a well-ordering). Another way to phrase this is that we are choosing the ‘first element of $X$ that showed up when we were building $L$.’
One thing to note is that, while $<_{L}$ is ‘canonical’ in the sense of being specifically definable, it is not ‘canonical’ in the sense of satisfying any more specific ‘naturality’ properties. If $\Lambda_{24}<_{L}\pi+e$, this doesn’t mean there’s any special relationship between the Leech lattice and the question of whether $\pi+e$ is an irrational number. Likewise there are arbitrary choices that go into the specific definition of $<_{L}$, so it could easily be the case that either $\ell^{2}<_{L}\mathbb{Q}_{19}$ or $\mathbb{Q}_{19}<_{L}\ell^{2}$ (depending on whether we like Hilbert spaces or the $p$-adic numbers more). Perhaps more pertinently to ordinary mathematical concerns, if we use $<_{L}$ to build objects one typically builds with the axiom of choice (e.g., maximal ideals, bases of vector spaces, and extensions of linear operators on Banach subspaces), there’s no reason to expect anything special about the resulting objects, aside from the fact that they’re explicitly definable.
For subsets of Polish spaces in particular, the definability of $<_{L}$ has the consequence that these kinds of choicy objects can be relatively simple to define in the sense of descriptive set theory. Specifically, Gödel showed [^7] Although in his notes, he credited the observation that the definition of the well-ordering can be phrased in such a simple way to Ulam [15, p. 197]. (See also [12, p. 151].) that in $L$ the projection of the complement of the projection of a Borel set may fail to be measurable.
@@
[…]
@@
Moreover, it can be arranged [^9] This is actually automatic for the definition of $<_{L}$ one typically writes down because every real number in $L$ is built at a countable stage of the transfinite construction of $L$. that $\{x\in\mathbb{R}:x<_{L}r\}$ is countable and $\{x\in\mathbb{R}:r<_{L}x\}$ is co-countable for every $r\in\mathbb{R}$, implying that $\pi_{3}\left(\mathbb{R}^{3}\setminus\pi_{4}(A)\right)$ is not Lebesgue measurable by Fubini’s theorem:
[…]
@@ 2. Effectively describing sets
Descriptive set theory as a field starts by asking the question ‘What can we say about regularity properties of subsets of $\mathbb{R}$ (and other metric spaces) that can be built using topologically tame sets, such as open and closed sets, using basic operations, such as (countable) unions and intersections, complements, and projections?’ It has its roots in the beginning of measure theory. Despite the fact that questions about measurability of simple-to-define sets are now largely settled, descriptive set theory continues to be an active area of research, with many applications in various areas of math (both within and outside of mathematical logic). It’s often useful in showing that classifying some family of objects up to isomorphism via ‘simply definable’ invariants is impossible.
@@
The most familiar examples of pointclasses are probably the classes of open and closed sets, but the classes of $G_{\delta}$ and $F_{\sigma}$ sets are also fairly common. Beyond $G_{\delta}$ and $F_{\sigma}$ is of course the pointclass of all Borel sets, and even further beyond this is the projective hierarchy (see Fig. 4), which is built out of Borel sets by successive applications of the operations of projections and complements and is a major topic of descriptive set theory and its interaction with set theory.
@@
Effective descriptive set theory grew out of a body of connections and analogies between descriptive set theory and computability theory. Within descriptive set theory more generally, it is often useful for fine-grained analysis and has applications to problems that do not overtly involve computability theory, such as work by Lutz, Qi, and Yu in [18] on the Hausdorff dimensions of Hamel bases.
@@
[…]
@@
The above exercise of choosing values for $N(X,n)$ gives a good example of how computability theorists (and other logicians for that matter) think. Most mathematicians engage with the natural numbers combinatorially or in an algebraic or analytic capacity. To a computability theorist, by contrast, a natural number is often just data. Any finite, possibly heterogeneous, bundle of data can be coded as a single natural number, as in our coding of $N(\mathbb{N}^{\mathbb{N}},n)$. Here we’re using $\nu_{2}$ in a fundamentally different way than the other $p$-adic valuations, in that it codes the length of initial segment of an element of $\mathbb{N}^{\mathbb{N}}$ that needs to be checked to verify membership in $N(\mathbb{N}^{\mathbb{N}},n)$. All of this is fairly unnatural from the point of view of number theory, but it doesn’t matter here. All that matters at the moment is that we can do it, just like with the definable well-order $<_{L}$ of $L$.
Recall that a set $W$ of natural numbers is computably enumerable if, informally speaking, there is a computer program that lists the elements of $W$ in some order. This can be formalized of course, but all reasonable formalizations of this notion result in an equivalent theory of computation. (This is the phenomenon of Turing completeness in computer programming.) $W$ is computable or decidable if there is a computer program that can decide membership of $W$. A basic exercise in computability theory is showing that $W$ is computable if and only if both $W$ and $\mathbb{N}\setminus W$ are computably enumerable. Famously Turing showed that there are computably enumerable sets that are not computable (specifically sets coding the solution to the halting problem), but there are also many examples in mathematics of sets that are known to be computably enumerable but not know to be computable, such as the set $\{|p-q|:p,q~{}\text{prime}\}$.
@@
Lusin and Sierpiński independently showed that one can build ‘universal $\undertilde{\mathbf{\Sigma}}^{1}_{n}$ sets.’ In other words, for any uncountable Polish space $X$, there is a $\undertilde{\mathbf{\Sigma}}^{1}_{n}$ set $A\subseteq X\times X$ such that for every $\undertilde{\mathbf{\Sigma}}^{1}_{n}$ set $W\subseteq X$, there is an $e\in X$ such that $\{x\in X:\langle x,e\rangle\in A\}=W$. This can be used to show that the (boldface and lightface) projective hierarchies are strict (see Fig. 4) via a diagonalization argument that is reminiscent of Turing’s proof of the unsolvability of the halting problem (which we can rephrase in our language here as the statement that there is a $\Sigma^{0}_{1}$ subset of $\mathbb{N}$ that is not $\Delta^{0}_{1}$).
@@
This follows from [12, Thm. 12.7] and 2.9. ∎

### v3

@@ Abstract.
is Lebesgue measurable. We also give similarly defined $g(x,y)$ with the property that the statement $x\mapsto g(x,r)$ is measurable for every $r\in\mathbb{R}$” has large cardinal consistency strength (and in particular implies the consistency of $\mathsf{ZFC}$) and $h(m,x,y,z)$ such that $h(1,x,y,z),\allowbreak\dots,\allowbreak h(16,x,y,z)$ can consistently be the indicator functions of a Banach–Tarski paradoxical decomposition of the sphere.
@@ Introduction
While most students of mathematics encounter measure theory at some point, the basic machinery of measure theory (such as $\sigma$-algebras) is often regarded as fussy and irrelevant in practice. In particular, the issue of measurability of sets and functions seems to be something one only encounters in the first semester of graduate real analysis and never again. The common informal explanation of this is that any function one can ‘actually write down’ will be measurable and that one needs to go looking for trouble using the axiom of choice in order to find non-measurable objects. This is certainly a true statement empirically and one would like it to be true on a moral level, but there is nevertheless a sense in which it is technically false and we will be exploring the extent to which this statement is both in some sense false and in some sense true here.
On a purely mathematical level, the main contributions of this paper are Theorems 4.8, 4.10, and 4.12. Each of these is a statement of the same form: “There is a degree $7$ polynomial $p$ with integer coefficients and fewer than $100$ variables such that writing some $\inf$’s and $\sup$’s over $\mathbb{N}$ and $\mathbb{R}$ in front of $p$ defines a function whose measurability is wildly independent of $\mathsf{ZFC}$.” More specifically we have the following:
  • (4.8) There is a degree $7$ polynomial $p(x,y,z,n,k_{1},\dots,k_{70})$ with integer coefficients such that it is independent of $\mathsf{ZFC}$ whether
    is Lebesgue measurable.
  • (4.10) There is a degree $7$ polynomial $p(x,y,z,w,t,n,k_{1},\dots,k_{74})$ with integer coefficients such that the function
    has the property that the statement “$g(x,r)$ is Lebesgue measurable for every $r\in\mathbb{R}$” has the same consistency strength as the existence of an inaccessible cardinal (i.e., a Grothendieck universe).
  • (4.12) There is a degree $7$ polynomial $p(m,x,y,z,w,t,n,k_{1},\dots,k_{76})$ with integer coefficients such that if we define
    $h(m,x,y,z)=\inf_{w\in\mathbb{R}}\sup_{t\in\mathbb{R}}\inf_{n\in\mathbb{N}}\sup_{\bar{k}\in\mathbb{N}^{76}}p(m,x,y,z,w,t,n,\bar{k}),$
    then it is independent of $\mathsf{ZFC}$ whether $h(1,x,y,z),\dots,h(16,x,y,z)$ are the indicator functions of a Banach–Tarski paradoxical decomposition of the unit sphere.
Although these results, to my [^1] We will occasionally use the terms I, me, and my as convenient shorthand for ‘the author’ or ‘the author’s,’ as appropriate. knowledge, have never been pointed out explicitly before, they are not morally new. 4.10 could have been written down in 1984 as a corollary of Shelah’s work showing that you ‘can’t take Solovay’s inaccessible away’ [29]. If we forget about the specific numbers of variables given, Theorems 4.8 and 4.12 could have been written down in 1970, when Matiyasevich (building on earlier work of Robinson, Davis, and Putnam) solved Hilbert’s tenth problem negatively [22]. If we moreover allow the polynomials to take some exponential functions as arguments (and if we only demand consistency with $\mathsf{ZFC}$ rather than independence), Theorems 4.8 and 4.12 could have been written down in 1960, when Robinson, Davis, and Putnam showed that every computably enumerable set can be realized as an exponential Diophantine set [7]. [^2] With a time machine and a copy of Green and Tao’s 2008 paper [11], this could be pushed back to 1959, when Davis and Putnam proved this under the assumption of the existence of arbitrarily long arithmetic sequences of prime numbers. If we allow ourselves to replace polynomials with arbitrary computable functions, these two results could have been written down at some point in the 50s, when the field of effective descriptive set theory was born out of the work of Kleene, Mostowski, and Addison. And finally if we just allow $p$ to be an arbitrary continuous function, then 4.8 was essentially written down in 1938 by Gödel in [10]. So it has been known for somewhere between 40 and 90 years that non-measurable objects can consistently be quite explicit, but despite how old these results are, it seems to me that not many mathematicians are aware of just how explicit these explicit objects can be.
Most mathematicians are aware of some of the more famous independence results, but aside from possibly a broad sketch of the proof of Gödel’s incompleteness theorem and maybe the word ‘forcing,’ very few know much about how you actually prove that you can’t prove something one way or another from some given axioms. Rather than just give the proofs of Theorems 4.8, 4.10, and 4.12, I’m going to try to tell the story of how one would even begin to prove such things, in part because the method of proof is entirely different from Gödelian self-reference and set-theoretic forcing, the two most well-known techniques for proving independence results, and in part because Theorems 4.8, 4.10, and 4.12 draw on several different important ideas from mathematical logic in the 20th century. Sections 1-4 of this paper essentially follow the events described in the previous paragraph in roughly chronological order. Section 1 discusses independence results generally and then describes Gödel’s inner model $L$, one of the core ideas in this paper and in set theory more broadly, in which one is able to well-order the reals in an ‘explicit’ way. Section 2 first deals with descriptive set theory, the study of the complexity of sets of real numbers that came out of early work in measure theory. Section 2 then discusses effective descriptive set theory, an extension of descriptive set theory that highlights analogies and connections between it and computability theory. Section 3 connects the ideas of the first two sections and in doing so allows us to upgrade our ‘explicit’ well-ordering of $\mathbb{R}$ to a ‘computable’ well-ordering of $\mathbb{R}$. Section 4 brings this all together with the famous negative resolution of Hilbert’s tenth problem in order to yield our main results, before discussing what would be involved in actually writing down the relevant polynomials (which we do not do).
The other side of the story—i.e., how it can be that objects like the above $f(x)$, $g(x,y)$, and $h(m,x,y,z)$ exist when in practice one seems to ‘need the axiom of choice’ to build non-measurable functions or sets—appears to some extent in the earlier sections but comes into focus in Section 5. A fair amount of work has been done by set theorists to explain this, such as Solovay’s famous construction (from an inaccessible cardinal) of a model of $\mathsf{ZF}+\mathsf{DC}$ in which all subsets of $\mathbb{R}$ are measurable [33]. In Section 5 we discuss this from the point of view of functions of a similar form to those appearing in Theorems 4.8, 4.10, and 4.12. Using work of Suslin and Lusin from classical descriptive set theory and work of van den Dries, Knight, Pillay, and Steinhorn on model-theoretic o-minimality, we show that many similarly defined functions are provably measurable in $\mathsf{ZFC}$. We then cite Shelah and Woodin’s proof that that ‘large cardinals imply that every reasonably definable set of reals is Lebesgue measurable’ in [30] to show that in the presence of sufficiently large large cardinals, anything like Theorems 4.8 and 4.12 becomes impossible in that the relevant functions are now provably measurable.
@@
Most of the notation we’ll use here is standard. To avoid collision with the notation for open intervals, we’ll use angle brackets for ordered tuples such as $\langle 1,2\rangle$. We will also be using the logician’s convention of representing tuples of indexed variables with a bar. For instance, $\bar{x}$ might represent $\langle x_{1},x_{2},x_{3}\rangle$. In some places these notations will be used together. For instance, if $\bar{x}=\langle x_{1},x_{2},x_{3}\rangle$, then $\langle\bar{x},n\rangle$ is the same thing as $\langle x_{1},x_{2},x_{3},n\rangle$. We’ll be representing projection maps with $\pi$, usually with a subscript. For example, an expression like $\pi_{4}$ is meant to represent a projection from some product $X\times Y\times Z\times W$ down to $X\times Y\times Z$, but this will be spelled out explicitly when it is used. Finally, we are taking 0 to be a natural number and all suprema and infima are understood to be computed in the extended real numbers, which we will denote by $\overline{\mathbb{R}}$.
@@ 1. Independence results and Gödel’s $L$
With the majority of set-theoretic independence phenomena, the very basic core argument is the same. One exhibits two models of something like $\mathsf{ZF}$ or $\mathsf{ZFC}$, a model satisfying a given statement and another model failing to satisfy that statement. The primary difference is that models of $\mathsf{ZFC}$ are large and rich enough to encode the majority of mathematics.
While models of set theory are obviously far harder to visualize than, say, groups with $6$ elements, sometimes the relative construction of these models can be visualized fairly clearly. Like many mathematicians, set theorists’ intuition is often founded on cartoonishly simple mental pictures, such as the standard drawing of a model of $\mathsf{ZFC}$ (see Figure 1). A model of $\mathsf{ZFC}$ is somewhat like a great tree, rooted at the empty set and with the ordinals forming a transfinitely tall trunk off of which the various sets coding familiar mathematical objects branch. This picture immediately gives two coordinates which describe the broad structure of the majority of constructions of one model of $\mathsf{ZF}(\mathsf{C})$ from another: One can imagine a model extending another in height, leading to the idea of large cardinals, or one can imagine a model that extends another in width, as is the case in set-theoretic forcing, or a model which is the result of pruning another to some narrower width, which is the very basic idea of inner models such as Gödel’s $L$. Many more elaborate proofs involve performing several (or even infinitely many) of these constructions in succession, such as in Cohen’s original proof of the independence of choice from $\mathsf{ZF}$, which involves forcing to widen a model and then pruning down to a narrower symmetric submodel (a kind of inner model specialized for the context of forcing) to kill the axiom of choice by carefully cutting certain choice functions out of the model.
@@ 1.1. The easiest independence result
For any inaccessible $\kappa\in M$, $H_{\kappa}$ is a model of $\mathsf{ZFC}$ (see Figure 2). This gives us what is probably the easiest $\mathsf{ZFC}$ independence result.
@@
Fix a model $M$ of $\mathsf{ZFC}$ with an inaccessible cardinal. Let $\kappa$ be the least inaccessible cardinal in $M$ (which exists by the fact that the cardinals are well-ordered). $H_{\kappa}^{M}$ is now a model of $\mathsf{ZFC}$ such that no set in $H_{\kappa}^{M}$ is itself a Grothendieck universe. Therefore $H_{\kappa}^{M}$ is a model of $\mathsf{ZFC}$ satisfying “there are no inaccessible cardinals.” Since we now have a model of $\mathsf{ZFC}$ with an inaccessible cardinal and a model of $\mathsf{ZFC}$ without an inaccessible cardinal, the statement “there is an inaccessible cardinal” must be independent of $\mathsf{ZFC}$. ∎
@@ 1.2. The narrowest inner model
Given a model $M$ of $\mathsf{ZF}$, a class $N\subseteq M$ is an inner model if $N$ contains all of the ordinals in $M$, NN is transitive (i.e., satisfies that if $x\in N$ and $y\in x$, then $y\in N$), and $N$ is a model of $\mathsf{ZF}$.
Gödel showed how to build the ‘narrowest’ inner model $L$, which can be thought of as the model of $\mathsf{ZF}$ ‘generated’ by the ordinals (see Figure 3), although this description is hard to make precise in a robust way. When it’s important to indicate that we are thinking about $L$ inside a particular model, we typically denote this with a superscript (i.e., $L^{M}$ is $L$ built in $M$).
@@ 1.3. The definable well-ordering of $L$
One particularly important aspect of the fine structure of $L$ is that there is a precise mechanism that results in the axiom of choice holding, unlike in arbitrary models of $\mathsf{ZFC}$. One way to understand this mechanism is that in the construction of $L$, sets are added ‘carefully’ so that it is possible to assign an ordinal ‘serial number’ to each set. In other words, there is a bijection $f:\operatorname{Ord}\to L$ that is definable in $L$. This induces a definable global well-ordering $_{L}$ of $L$, which then gives a definable global choice function: Whenever you need to pick an element of a set $X$, you can just take the $_{L}-$least element of $X$ (which always exists since $_{L}$ is a well-ordering). Another way to phrase this is that we are choosing the ‘first element of $X$ that showed up when we were building $L$.’
One thing to note is that, while $_{L}$ is ‘canonical’ in the sense of being specifically definable, it is not ‘canonical’ in the sense of satisfying any more specific ‘naturality’ properties. If $\Lambda_{24}<_{L}\pi+e$, this doesn’t mean there’s any special relationship between the Leech lattice and the question of whether $\pi+e$ is an irrational number. Likewise there are arbitrary choices that go into the specific definition of $_{L}$, so it could easily be the case that either $\ell^{2}<_{L}\mathbb{Q}_{19}$ or $\mathbb{Q}_{19}<_{L}\ell^{2}$ (depending on whether we like Hilbert spaces or the $p$-adic numbers more). Perhaps more pertinently to ordinary mathematical concerns, if we use $_{L}$ to build objects one typically builds with the axiom of choice (e.g., maximal ideals, bases of vector spaces, and extensions of linear operators on Banach subspaces), there’s no reason to expect anything special about the resulting objects, aside from the fact that they’re explicitly definable.
For subsets of Polish spaces in particular, the definability of $_{L}$ has the consequence that these kinds of choicy objects can be relatively simple to define in the sense of descriptive set theory. Specifically, Gödel showed [^7] Although in his notes, he credited the observation that the definition of the well-ordering can be phrased in such a simple way to Ulam [18, p. 197]. (See also [15, p. 151].) that in $L$ the projection of the complement of the projection of a Borel set may fail to be measurable.
@@
Assuming V=LV=L, [^8] V=LV=L is the set-theoretic axiom that states ‘all sets are in $L$.’ $V$ is the notation for the class of all sets in a model of $\mathsf{ZF}$. In other words, $M$ satisfies $V=L$ if and only if $M=L^{M}$. there is a $G_{\delta}$ set (i.e., a countable intersection of open sets) $A\subseteq\mathbb{R}^{4}$ such that
@@
Moreover, it can be arranged [^9] This is actually automatic for the definition of $_{L}$ one typically writes down because every real number in $L$ is built at a countable stage of the transfinite construction of $L$. that $\{x\in\mathbb{R}:x<_{L}r\}$ is countable and $\{x\in\mathbb{R}:r<_{L}x\}$ is co-countable for every $r\in\mathbb{R}$, implying that $\pi_{3}\left(\mathbb{R}^{3}\setminus\pi_{4}(A)\right)$ is not Lebesgue measurable by Fubini’s theorem:
$\int_{0}^{1}\int_{0}^{1}\begin{rcases}\begin{dcases}1&x<_{L}y\\ 0&x\not<_{L}y\end{dcases}\end{rcases}dxdy=0\neq 1=\int_{0}^{1}\int_{0}^{1}\begin{rcases}\begin{dcases}1&x<_{L}y\\ 0&x\not<_{L}y\end{dcases}\end{rcases}dydx.$
@@ 2. Effectively describing sets
Descriptive set theory as a field starts by asking the question ‘What can we say about regularity properties of subsets of $\mathbb{R}$ (and other metric spaces) that can be built using topologically tame sets, such as open and closed sets, using basic operations, such as (countable) unions and intersections, complements, and projections?’ It has its roots in the beginning of measure theory and the work of Cantor, Borel, Lebesgue, Suslin, Lusin, and Sierpiński. Despite the fact that questions about measurability of simple-to-define sets are now largely settled, descriptive set theory continues to be an active area of research, with many applications in various areas of math (both within and outside of mathematical logic). It’s often useful in showing that classifying some family of objects up to isomorphism via ‘simply definable’ invariants is impossible.
@@
The most familiar examples of pointclasses are probably the classes of open and closed sets, but the classes of $G_{\delta}$ and $F_{\sigma}$ sets are also fairly common. Beyond $G_{\delta}$ and $F_{\sigma}$ is of course the pointclass of all Borel sets, and even further beyond this is the projective hierarchy (see Figure 4), which is built out of Borel sets by successive applications of the operations of projections and complements and is a major topic of descriptive set theory and its interaction with set theory.
@@
$\undertilde{\mathbf{\Sigma}}^{1}_{1}$$\undertilde{\mathbf{\Sigma}}^{1}_{2}$$\undertilde{\mathbf{\Sigma}}^{1}_{3}$$\text{Borel}=\undertilde{\mathbf{\Delta}}^{1}_{1}$$\undertilde{\mathbf{\Delta}}^{1}_{2}$$\undertilde{\mathbf{\Delta}}^{1}_{3}$$\cdots$$\undertilde{\mathbf{\Pi}}^{1}_{1}$$\undertilde{\mathbf{\Pi}}^{1}_{2}$$\undertilde{\mathbf{\Pi}}^{1}_{3}$$\subsetneq$$\subsetneq$$\cdots$$\subsetneq$$\subsetneq$$\subsetneq$$\subsetneq$$\subsetneq$$\subsetneq$$\subsetneq$$\subsetneq$$\cdots$}
Effective descriptive set theory grew out of a body of connections and analogies between descriptive set theory and computability theory and specifically the work of Kleene, Mostowski, and Addison in the 1950s. Within descriptive set theory more generally, it is often useful for fine-grained analysis and has applications to problems that do not overtly involve computability theory, such as work by Lutz, Qi, and Yu in [21] on the Hausdorff dimensions of Hamel bases.
@@
A basic Polish space is a finite product of the spaces $\mathbb{N}$, $\mathbb{N}^{\mathbb{N}}$, $\mathbb{R}$, and $\overline{\mathbb{R}}$ (where $\overline{\mathbb{R}}\coloneq[-\infty,\infty]$ is the extended real numbers).
@@
The above exercise of choosing values for $N(X,n)$ gives a good example of how computability theorists (and other logicians for that matter) think. Most mathematicians engage with the natural numbers combinatorially or in an algebraic or analytic capacity. To a computability theorist, by contrast, a natural number is often just data. Any finite, possibly heterogeneous, bundle of data can be coded as a single natural number, as in our coding of $N(\mathbb{N}^{\mathbb{N}},n)$. Here we’re using $\nu_{2}$ in a fundamentally different way than the other $p$-adic valuations, in that it codes the length of initial segment of an element of $\mathbb{N}^{\mathbb{N}}$ that needs to be checked to verify membership in $N(\mathbb{N}^{\mathbb{N}},n)$. All of this is fairly unnatural from the point of view of number theory, but it doesn’t matter here. All that matters at the moment is that we can do it, just like with the definable well-order $_{L}$ of $L$.
Recall that a set $W$ of natural numbers is computably enumerable if, informally speaking, there is a computer program that lists the elements of $W$ in some order. This can be formalized of course, but all reasonable formalizations of this notion result in an equivalent theory of computation. (This is the phenomenon of Turing completeness in computer programming.) $W$ is computable or decidable if there is a computer program that can decide membership of $W$. A basic exercise in computability theory is showing that $W$ is computable if and only if both $W$ and $\mathbb{N}\setminus W$ are computably enumerable. Famously Turing showed that there are computably enumerable sets that are not computable (specifically sets coding the solution to the halting problem), but there are also many examples in mathematics of sets that are known to be computably enumerable but not known to be computable, such as the set $\{|p-q|:p,q~\text{prime}\}$.
@@
Lusin and Sierpiński independently showed that one can build ‘universal $\undertilde{\mathbf{\Sigma}}^{1}_{n}$ sets.’ In other words, for any uncountable Polish space $X$, there is a $\undertilde{\mathbf{\Sigma}}^{1}_{n}$ set $A\subseteq X\times X$ such that for every $\undertilde{\mathbf{\Sigma}}^{1}_{n}$ set $W\subseteq X$, there is an $e\in X$ such that $\{x\in X:\langle x,e\rangle\in A\}=W$. This can be used to show that the (boldface and lightface) projective hierarchies are strict (see Figure 4) via a diagonalization argument that is reminiscent of Turing’s proof of the unsolvability of the halting problem (which we can rephrase in our language here as the statement that there is a $\Sigma^{0}_{1}$ subset of $\mathbb{N}$ that is not $\Delta^{0}_{1}$).
@@
This follows from [15, Thm. 12.7] and 2.9. ∎